OpenStudy (anonymous):

If 4(2x + 2) - (2x+ 5) = -3 then x = A. 1 B. -1 C. -8/3 D. 4

OpenStudy (blank ):

x = -1

OpenStudy (jadeishere):

So, the answer would be B, x = -1 here's why; Reorder the terms -2x + 4(2x + 2) - 5 = -3 Distribute 4 to 2x and 2; -2x + 8x + 8 - 5 = -3 Combine like terms; 6x + 3 = -3 Isolate the variable; 6x = -6 Complete; x = -1 Hope this helps :)